AddieroniAndCheese:

GUYS ITS DUE IN 2 HOURS AND I CANT DO HALF REACTIONS HELLLLP PLEASE ap chem a) Identify the oxidation and reduction half-reactions for the equation below and describe the number of electrons that are involved in the reaction. 2Ag+(aq) + Pb(s) → 2Ag(s) + Pb2+(aq) b) Explain the reaction that would occur in a galvanic cell between solutions of Al3+ and Pb2+. Identify the oxidation and reduction reactions and the location of each reaction. Then, use the standard reduction potential tables to determine the cell potential.

AddieroniAndCheese:

1 attachment
AddieroniAndCheese:

1 attachment
gdubb08:

To analyze this redox reaction, let's break it down into its oxidation and reduction half-reactions and determine the number of electrons involved. 1. Identifying the half-reactions: Reduction half-reaction: Ag+ (aq) + e- → Ag(s) Oxidation half-reaction: Pb(s) → Pb2+ (aq) + 2e- 2. Number of electrons involved: In the reduction half-reaction: - Each Ag+ ion gains one electron to become Ag(s) - There are two Ag+ ions in the overall equation - Therefore, 2 electrons are gained in total In the oxidation half-reaction: - Each Pb atom loses two electrons to become Pb2+ - There is one Pb atom in the overall equation - Therefore, 2 electrons are lost in total The number of electrons lost in the oxidation half-reaction equals the number of electrons gained in the reduction half-reaction, which is 2 electrons. In summary: - Reduction: 2Ag+ (aq) + 2e- → 2Ag(s) - Oxidation: Pb(s) → Pb2+ (aq) + 2e- - Total number of electrons transferred: 2 electrons

gdubb08:

In a galvanic cell between solutions of Al3+ and Pb2+, the following reaction would occur: 2Al(s) + 3Pb2+(aq) → 2Al3+(aq) + 3Pb(s) The half-reactions and their locations are: Oxidation (occurs at the anode): Al(s) → Al3+(aq) + 3e- Reduction (occurs at the cathode): Pb2+(aq) + 2e- → Pb(s) To determine the cell potential, we need to use the standard reduction potentials: Al3+(aq) + 3e- → Al(s) E° = -1.66 V Pb2+(aq) + 2e- → Pb(s) E° = -0.13 V The cell potential (E°cell) is calculated by subtracting the oxidation potential from the reduction potential: E°cell = E°reduction - E°oxidation E°cell = E°(Pb2+/Pb) - E°(Al3+/Al) E°cell = (-0.13 V) - (-1.66 V) = 1.53 V Therefore, the standard cell potential for this galvanic cell is 1.53 V[1][2].

AddieroniAndCheese:

TYSM!

gdubb08:

@addieroniandcheese wrote:
TYSM!
np